# Histogram Layers for Neural “Engineered” Features

### v1

@@ V-B Ablation Study: Multichannel processing
The results of our multichannel processing for the best NLBP and NEHD from Section V-A are shown in Table III. From the results, independently processing each channel achieved the highest average test accuracy. Independently processing each channel leads to the most information so this result is intuitive. However, the number of features will scale as the number of input feature channels are increased. When observing the results of converting to gray scale and applying a $1\times 1$ convolution, the average test accuracy is comparable to the independent processing approach, though these approaches have increased variability indicated by the larger standard deviation when compared to the standard deviation of the independent processing approach. When integrating the neural “engineered” features into a deeper network, the $1\times 1$ approach may be the most ideal approach as this can reduce the number of features extracted when compared to the independent fusion approach. The grayscale conversion is limited to RGB inputs.
V-C Comparison of Neural and “Engineered” Features
@@
The last set of experiments compared our proposed NEHD and NLBP with the baseline EHD and LBP variants across three different datasets: FashionMNIST [xiao2017fashion], PRMI [xu2022prmi], BloodMNIST [medmnistv1, medmnistv2, bloodmnist]. The results of the comparisons are shown in Table IV. Overall, both NEHD and NLBP outperform the “engineered” features across all three datasets. For FashionMNIST, as discussed in Section V-A, edges are important to distinguish the different articles of clothes resulting in EHD achieving the best performance for “engineered” features. However, our NEHD and NLBP achieve statistically significantly (i.e., no overlapping error bars) improved performance. PRMI and BloodMNIST are consists of RGB images. Each feature was applied to each channel independently as a fair comparison between each approach. The PRMI results show that NLBP slightly outperformed NEHD. For the PRMI datasets, edge information is important to identify the roots, but there are some images have some illumination differences. LBP is robust to illumination variations (e.g., monotonic changes) [humeau2019texture], resulting in NLBP retaining a similar property when extracting texture features from the images.
@@
The NEHD and NLBP are compared with the best EHD and LBP variant qualitatively (Figure 6) and quantitatively (Figure 7). As seen in the TSNE visuals in Figure 6, the features learned by NEHD and NLBP appear more compact in the 2D projection than the “engineered” features. An important aspect of both approaches is the aggregation performed by the histogram layer. The soft binning approach provides a way to account for intra-class variations that can be present within the dataset [peeples2021histogram]. As noted in the confusion matrices in Figure 7, the neural “engineered” features outperform the “engineered” features. For some of the PRMI images, the root are occluded with background information (e.g., soil, image artifacts) and the fixed representation of the “engineered” features is limited in mitigating the impact of these confusers in the images. However, our neural “engineered” features can change the texture representation to improve classification performance.
@@
BloodMNIST was the most difficult dataset for the shallow network used since the number of classes was eight (more than PRMI) in comparison and the blood cells visually have small differences between the different classes. However, our proposed features achieved statistically significant higher test accuracy than the “engineered” features. Texture information is vital in several biomedical applications [liu2019bow]. Our proposed approach can be applied to several biomedical tasks and integrated into deeper networks to achieve improved peformance on the BloodMNIST dataset.
VI Conclusion
@@
In this work, we proposed neural handcrafted features by using histogram layers to aggregate structural texture features. We demonstrate the approach by introducing two neural handcrafted features: NEHD and NLBP. Our results across benchmark (FashionMNIST) and real-world (PRMI and BloodMNIST) datasets showcase the potential use of these features. The general framework for neural handcrafted features can be used for other texture feature approaches such as Haralick texture features [haralick1973textural], histogram of oriented gradients [dalal2005histograms] and several more statistical and structural texture feature extraction methods [liu2019bow, humeau2019texture]. Future work includes integrating the neural handcrafted layer(s) into deeper networks, improving the multichannel processing approach, and designing new objective functions to maximize statistical and/or structural texture information.
@@ Acknowledgment
[…]
[Uncaptioned image] Joshua Peeples received the Ph.D. degree in electrical and computer engineering from the University of Florida in 2022. He is currently an Assistant Professor in the Department of Electrical and Computer Engineering at Texas A&M University. His primary research interests include machine learning, computer vision, and image processing with a focus on image texture analysis.
[Uncaptioned image] Salim Al Kharsa is an undergraduate in the Department of Electrical and Computer Engineering at Texas A&M University. He has an expected graduation date of May 2024.
[Uncaptioned image] Luke Saleh is an undergraduate Computer Engineering student at the University of Florida with an expected graduation date of December 2024. He is currently working as an undergraduate researcher for the Machine Learning and Sensing Lab at UF. His primary interests include embedded systems, machine learning, digital hardware design, and data science.
[Uncaptioned image] Alina Zare Alina Zare (Senior Member, IEEE) teaches and conducts research in the area of machine learning and artificial intelligence as a Professor in the Electrical and Computer Engineering Department at the University of Florida. She also serves at the Associate Dean for Research and Facilities at the Herbert Wertheim College of Engineering. Dr. Zare’s research has focused primarily on developing new machine learning algorithms to automatically understand and process data and imagery. Her research work has included automated plant root phenotyping, sub-pixel hyperspectral image analysis, target detection and underwater scene understanding using synthetic aperture sonar, LIDAR data analysis, Ground Penetrating Radar analysis, and buried landmine and explosive hazard detection.

### v2

@@ 5.4 Comparison of Neural and “Engineered” Features
The last set of experiments compared our proposed NEHD and NLBP with the baseline EHD, LBP variants, and lightweight CNNs across three different datasets: FashionMNIST [50], PRMI [52], BloodMNIST [53, 54, 55]. The results of the comparisons are shown in Table 5. Overall, both NEHD and NLBP outperform the “engineered” features across all three datasets. For FashionMNIST, as discussed in Section 5.1, edges are important to distinguish the different articles of clothing resulting in EHD achieving the best performance for “engineered” features. However, our NEHD and NLBP achieve statistically significantly (i.e., no overlapping error bars) improved performance. When compared with the DSANet [56] and MSDCNN [57], our proposed NEHD outperforms these two models on this datasets and NLBP performs better than MSDCNN as well. This shows the power of these learnable histogram-based texture features can potentially capture useful information more directly than deeper CNNs. Also, our proposed models have significantly less learnable parameters than the lightweight CNNs. For FashionMNIST, DSANet and MSDCNN had a total of 290,894 and 630,104 respectively. The proposed NEHD and NLBP have 43,732 and 92,282 respectively which further supports the potential of these neural “engineered” features. If only the feature extraction layers learnable parameters are considered, NEHD only has 162 while NLBP has 112.
@@
PRMI and BloodMNIST are consists of RGB images. Each feature was applied to each channel independently as a fair comparison between each approach. The PRMI results show that NLBP slightly outperformed NEHD. For the PRMI dataset, edge information is important to identify the roots, but there are some images have some illumination differences. LBP is robust to illumination variations (e.g., monotonic changes) [20], resulting in NLBP retaining a similar property when extracting texture features from the images. NEHD and NLBP also performed well on the PRMI dataset in comparison to the DSANet and MSDCNN. The difference in performance of the MSDCNN vs DSANet may be due to the multi-scale convolutional approach employed in the network. However, the results on the PRMI further support the usefulness of texture features especially in Agriculture focused applications [63].
@@
The NEHD and NLBP are compared with the best EHD and LBP variant qualitatively (Figure 7) and quantitatively (Figure 8). As seen in the TSNE visuals in Figure 7, the features learned by NEHD and NLBP appear more compact in the 2D projection than the “engineered” features. An important aspect of both approaches is the aggregation performed by the histogram layer. The soft binning approach provides a way to account for intra-class variations that can be present within the dataset [10]. As noted in the confusion matrices in Figure 8, the neural “engineered” features outperform the “engineered” features. For some of the PRMI images, the roots are occluded with background information (e.g., soil, image artifacts) and the fixed representation of the “engineered” features is limited in mitigating the impact of these confusers in the images. However, our neural “engineered” features can change the texture representation to improve classification performance.
@@
BloodMNIST was the most difficult dataset for the shallow networks used since the number of classes was eight (more than PRMI) in comparison and the blood cells visually have small differences between the different classes. However, our proposed features achieved statistically significant higher test accuracy than the “engineered” features. Texture information is vital in several biomedical applications [8]. Our proposed approach can be applied to several biomedical tasks and integrated into deeper networks to achieve improved performance on the BloodMNIST dataset. The lightweight CNNs did perform better on this dataset than NEHD and NLBP. As discussed, this was the most challenging dataset of the three and deeper networks will improve the feature representations learned for classifying the blood cells. An interesting observation is that the multi-scale network (MSDCNN) performed the best on this datasets. Future studies could investigate adapting NEHD and NLBP for multi-scale texture analysis especially since texture is very scale dependent [63, 64].
6 Conclusion
@@
In this work, we proposed neural “engineered” features by using histogram layers to aggregate structural texture features. We demonstrate the approach by introducing two neural “engineered” features: NEHD and NLBP. Our results across benchmark (FashionMNIST) and real-world (PRMI and BloodMNIST) datasets showcase the potential use of these features. The general framework for neural “engineered” features can be used for other texture feature approaches such as Haralick texture features [15], histogram of oriented gradients [13] and several more statistical and structural texture feature extraction methods [8, 20]. Future work includes integrating the neural handcrafted layer(s) into deeper networks, improving the multichannel processing approach, and designing new objective functions to maximize statistical and/or structural texture information.